I need this quick please

27. A pure solvent freezes at 43.4°C. The freezing temperature is 40.7°C when 52.5 g of CO(NH-h is in 1.00 x 102 g of the solvent. Calculate the molal freezing point constant of the solvent. (The molar mass for CO(NH2h is 60.06 g/mol.) a. 0.309 °C/m b. 0.343°C/m c. 0.93°C/m d. 13.9°C/m e. 14.9°C/m 28. K-1.3 x 10-10 for the reaction HC,H,O(a)H (aq) + C,H,O (aq). In which direction is the equilibrium favored? a. Forward b. Reverse c. Appreciable quantities of all species are present at equilibrium d. No direction is favored e. There is not enough information to determine the direction
Answer #1

27. When freezing point decreases after the addition of solute this is called depression in freezing point and can be calculated by using the formula:

?Tf = Kf ×m×i

Kf = ?Tf/(m×i)

?Tf = Tf(solution) - T°f(pure solvent)=(43.4-40.7)?=2.7?

m= molality = moles of solute/mass of solvent in kg

Moles of solute= given mass/it's molar mass

= 52.5g/60.06g/mol =0.874 mol

Mass of solvent= 1.00×10^2g = 0.1 kg

So, m= 0.874 mol/0.1kg = 8.74 m

For non dissociative or non associative solute i =1

So, putting all the values in the formula

Kf = 2.7?/(8.74 m)=0.309?/m

So the correct option is a)

28. At equilibrium the rate of forward reaction is equal to the rate of backward reaction.

And the equilibrium constant K gives the information about the ratio of concentration of products to the concentration of reactants.

This means equilibrium constant donot give any information about the direction of reaction. So,

The correct option is e)

* Its the reaction quotient(Q) that can give the idea about the direction of the reaction.

If Q>K then reaction will proceed in backward direction.

If Q<K then the reaction will proceed in forward direction